Suppose v1v2..vmis linearly independent in V and w element of V. Show that v1,v2,...vm;w is linearly independent if and only if w = not element of span(v1,v2,...vm).


1
Expert's answer
2021-05-17T18:56:46-0400

Let "v_1,v_2,...,v_m," w is linearly dependent and let

"w\in span(v_1,v_2,...v_m)"


"w=\alpha_1v_1+\alpha_2v_2+....+\alpha_mv_m=0," Where at least one of "\alpha_i\neq 0."


"\Rightarrow \alpha_1v_1+\alpha_2v_2+....+\alpha_mv_m-w=0"


"\Rightarrow \alpha_1v_1+\alpha_2v_2+....+\alpha_mv_m+\alpha_{m+1}w=0,\text{ Where }\alpha_{m+1}=-1\neq 0"


So, ("v_1,v_2,...v_m" ), w are linearly independent which is contradiction.

Hence, "w\in span((v_1,v_2,...v_m)"


Conversely: Let "w \in span(v_1,v_2,..,v_m)"


TO show: "x_1,x_2...,x_m , w" is linearly independent

then, "\alpha_1,\alpha_2,.....,\alpha_{m+1}\neq 0"


Such that,

 "\alpha_1v_1+\alpha_2v_2+....+\alpha_mv_m+\alpha{m+1}w=0"


 say "\alpha_{m+1}\neq 0."


"\alpha_{m+1} w=-\alpha_1v_1-\alpha_2v_2.... -\alpha_mv_m"


"w=-\alpha_1\alpha^{-1}_{m+1}v_1-\alpha_2\alpha^{-1}_{m+1}2v_2.... -\alpha_m\alpha^{-1}_{m+1}v_m"


"w\in span(v_1,v_2,...,v_m)," Which is contradiction.


Hence , "v_1,v_2,...,v_m , w" is linearly independent.